# Interpretability by Design for Efficient Multi-Objective Reinforcement Learning

### v1

@@ 3.5 The LLE-MORL Algorithm
The locally linear extension process is critical for the LLE-MORL algorithm in tracing an approximate Pareto front efficiently. The full algorithm (see also App. A) consists of five stages: (1) Initialization: We train a set of $K$ base policies $\{\theta_{w_{i}}\}_{i=1}^{K}$ to convergence using PPO (Schulman et al., 2017). Each policy is trained under a distinct scalarization weight $w_{i}\in\Omega$, where these weights are chosen to be evenly distributed across the preference space. (2) Directional Retraining: For each $i=1,\dots,K-1$, continue train based on $\theta_{w_{i}}$ under a new preference $w_{i^{\prime}}$ for $T_{\mathrm{dir}}$ steps to obtain $\theta_{w_{i}^{\prime}}$, where $\theta_{w_{i}}$ and $\theta_{w_{i^{\prime}}}$ should be both non-dominated points. Record the parameter update vector $\Delta\theta_{i}=\theta_{w_{i}^{\prime}}-\theta_{w_{i}}$ and weight shift $\Delta w_{i}=w_{i+1}-w_{i}$. (3) Locally Linear Extension: For each base policy $\theta_{w_{i}}$, we generate a set of intermediate policies by applying each step-scale factor $\alpha_{j}$ to the parameter update vector $\Delta\theta_{i}$. Concretely, each candidate is $\theta_{i,j}=\theta_{w_{i}}+\alpha_{j}\,\Delta\theta_{i}$, allowing negative and positive moves along the local direction in parameter space. Simultaneously, we adjust the preference weight by $\Delta w_{i}$ scaled by $\alpha_{j}$ to obtain $w_{i,j}$. These step-scale factors control how far along the local direction each intermediate policy moves. (4) Candidate Selection: All candidate policies $\theta_{i,j}$ generated in the locally linear extension stage are evaluated to obtain their respective performance vectors. From this set of extended policies, we identify and select the subset of non-dominated solutions. These selected non-dominated candidates are then advanced to the fine-tuning stage. (5) Preference-Aligned Fine-Tuning: from each candidate $\theta$ and its matched weight $w$, perform a short PPO fine-tuning of $T_{\mathrm{ref}}$ steps under $w$ to push the generated policy closer to the true Pareto front.
@@ 4 Experiments
4.1 Benchmark Problems
In this section, we evaluate the LLE-MORL algorithm using popular continuous MORL benchmark problems from the MO-Gymnasium (Felten et al., 2023). Our benchmark problems include three two-objective continuous environments.
MO-Swimmer-v5. A planar, three-link swimmer operating in a viscous fluid, utilising a 2D continuous action space to control its joint torques. The objectives are to maximise forward velocity along the $x$-axis and minimise the control cost.
MO-Hopper-2d-v5. This environment features a 2D one-legged hopper with a 3-dimensional continuous action space controlling torques for its thigh, leg, and foot joints. Originally a 3-objective task (forward speed, jump height, control cost), we use the 2-objective variant, in which the separate control-cost objective is added to other objectives.
MO-Ant-2d-v5. A quadrupedal “ant” robot in 2D with an eight-dimensional action space for joint torques. By default, the environment emits a three-dimensional reward vector: (1) $x$-velocity, (2) $y$-velocity, and (3) control cost. Here, we use the two-objective variant in which the separate control-cost objective is added to other objectives.
4.2 Evaluation Metrics
@@
4.3 Baselines
We compare our LLE-MORL against the following state-of-the-art MORL algorithms: (i) GPI-LS (Alegre et al., 2023) applies Generalised Policy Improvement over a discretised set of preference weights and uses linear scalarization to construct a diverse Pareto set. (ii) A Conservative Approximate Pareto Q-Learning method (CAPQL) (Lu et al., 2023): learns an ensemble of Q-functions under different preferences and selects actions via conservative aggregation to improve front coverage. (iii) MORL/D (Felten et al., 2024) is a deep-RL analogue of decomposition-based multi-objective optimisation that trains subpolicies under scalarised objectives and recombines them via weight decompositions to approximate the Pareto front.
4.4 Results and Analysis
(a) MO-Swimmer
(b) MO-Hopper-2d
(c) MO-Ant-2d
Figure 4: Pareto fronts from the sample-efficient training setting, comparing our LLE-MORL method with baselines on three continuous-control benchmarks. LLE-MORL demonstrates more comprehensive Pareto fronts across all benchmarks.
(a) MO-Swimmer
(b) MO-Hopper-2d
(c) MO-Ant-2d
Figure 5: Pareto fronts from the standard-training setting, comparing our LLE-MORL method with baselines on three continuous-control benchmarks. LLE-MORL consistently achieves wider coverage and closer proximity to the true Pareto front.
@@ 4.4 Results and Analysis
To assess the performance of MORL, we now present quantitative results evaluating the quality of the approximated Pareto fronts. We conduct experiments under two distinct settings to provide a comprehensive understanding of algorithms’ capabilities: (1) Sample-Efficient Setting: All methods, including our LLE-MORL approach, were trained for $1.5\times 10^{5}$ timesteps. Given the complexity of continuous control benchmarks, this relatively limited interaction budget serves as a critical testbed for evaluating how rapidly different MORL strategies can discover effective Pareto front approximations. (2) Standard-Training Setting: To assess performance under more common training conditions for these continuous control benchmarks, most methods, including our LLE-MORL approach, were trained for $1\times 10^{6}$ timesteps. An exception was made for the CAPQL baseline, which, due to its significant computational demands, was trained for $5\times 10^{5}$ timesteps. This setting aligns with common practices for benchmarking in continuous control and allows us to assess the final quality of the Pareto fronts achieved by each algorithm after a more thorough learning process. Detailed training setups can be found in App. B.
@@
First, we analyse performance in the sample-efficient setting, with results using Hypervolume (HV), Expected Utility (EU), and Sparsity (SP) metrics presented in Table 1 and the corresponding Pareto front visualisations in Figure 4. In this limited-interaction scenario, LLE-MORL achieves the highest HV and EU in all benchmarks, demonstrating strong capabilities in rapidly achieving high-quality Pareto fronts. Regarding SP, while LLE-MORL does not consistently achieve the leading scores on this metric, its performance generally reflects a good and effective distribution of solutions along the high-quality Pareto fronts it identifies. It should be noted that SP results can be confounded by a fragmentary recovery of the Pareto front. For instance, if CAPQL discovers only two close points of the Pareto front for the MO-Ant problem, the sparsity rating is nearly perfect. Conversely, a low SP might also arise from solutions being overly clustered in a small region, as potentially seen with GPI-LS in MO-Swimmer shown in Figure 4(a).
@@
Transitioning to the standard-training setting, where methods were trained for a more extensive duration, the evaluation results are presented in Table 2, and the corresponding Pareto front visualisation can be found in Figure 5. Across all benchmarks, LLE-MORL typically achieves the highest HV and highly competitive EU. This superior performance indicates that LLE-MORL finds a more extensive and higher-quality set of solutions, which strongly suggests a better approximation of the Pareto front compared to the baselines. The evidence from Pareto front visualisation further corroborates LLE-MORL’s advantages. In the MO-Swimmer environment, shown in Figure 5(a), LLE-MORL more comprehensively explores the objective space, successfully identifying Pareto optimal solutions in the lower-right region consistently missed by baselines such as GPI-LS and CAPQL. Notably, when comparing the GPI-LS and CAPQL performance to those in the sample-efficient setting for the MO-Swimmer environment, these particular baselines appear to remain constrained by suboptimal solutions in this challenging region, indicating that simply extending training duration did not resolve their exploration deficiencies here. While LLE-MORL’s thorough exploration to achieve this broader coverage means its SP may not be the numerically lowest, the result could be well-justified by the extensive nature of the front.
In summary, LLE-MORL consistently demonstrates superior Pareto front approximations across both sample-efficient and standard-training evaluations. This robust performance is significantly supported by its innovative extension process, which is largely training-free once core parameter-performance relationships are established, allowing for the efficient generation of diverse and high-quality solutions. Consequently, LLE-MORL excels at both rapid learning in data-limited scenarios and achieving comprehensive, high-fidelity fronts with extended training, highlighting its distinct advantages for multi-objective reinforcement learning. Additionally, we present the running time of each algorithm in App. B, which shows the high efficiency level achievable by LLE-MORL.
@@
GPI-LS CAPQL MORL/D LLE-MORL-0 LLE-MORL
MO-Swimmer HV$(10^{4})$ 4.92 5.16 5.95 6.68 6.93
EU$(10^{1})$ 0.99 1.09 1.01 0.84 1.09
SP$(10^{2})$ 0.05 0.97 21.99 0.88 0.66
MO-Hopper HV$(10^{5})$ 1.22 0.95 1.96 2.68 2.77
EU$(10^{2})$ 2.31 2.02 3.33 4.05 4.13
SP$(10^{2})$ 0.57 4.48 43.10 16.56 33.85
MO-Ant HV$(10^{4})$ 2.22 8.52 9.65 10.44 10.67
EU$(10^{2})$ 0.46 1.91 1.99 2.25 2.32
SP$(10^{3})$ 2.46 0.27 1.11 1.45 1.14
Table 1: Sample-efficient evaluation of the quality of the Pareto front by hypervolume (HV), expected utility (EU) and sparsity (SP).
Environment Metric Method
GPI-LS CAPQL MORL/D LLE-MORL-0 LLE-MORL
MO-Swimmer HV$(10^{4})$ 5.56 4.86 6.68 7.37 7.44
EU$(10^{1})$ 1.09 1.10 1.10 1.03 1.03
SP$(10^{2})$ 0.07 0.02 6.38 2.10 2.28
MO-Hopper HV$(10^{5})$ 1.13 1.70 3.62 4.88 5.05
EU$(10^{2})$ 2.26 3.14 5.02 5.65 5.77
SP$(10^{2})$ 1.91 35.94 4.93 7.20 6.56
MO-Ant HV$(10^{5})$ 2.44 1.62 1.28 2.43 2.51
EU$(10^{2})$ 3.55 2.93 2.56 3.44 3.61
SP$(10^{3})$ 0.44 0.06 9.19 1.76 1.33
Table 2: Standard-training evaluation of the quality of Pareto front by hypervolume (HV), expected utility (EU) and sparsity (SP).

### v2

@@ 3.4 The LLE-MORL Algorithm
Motivated by empirical observations, LLE-MORL is built on the hypothesis that there exists a manifold in parameter space that corresponds locally to the Pareto front and has the same intrinsic dimension, along which multi-objective performance varies smoothly. Theoretical analysis and formal proofs are provided in Section 3.5 and Appendix B.
@@
Consider a $d$-objective MORL problem. The full Algorithm 1 (see Appendix A) consists of five stages: (1) Initialisation: We train a set of $K$ base policies $\{\theta_{w_{k}}\}_{k=1}^{K}$ to convergence using PPO (Schulman et al., 2017). Each policy is trained under a distinct scalarization weight $w_{k}\in\Omega$, where these weights are chosen to be evenly distributed across the preference space. (2) Directional Retraining: For each base policy $\theta_{w_{k}}(k=1,\dots,K)$, we estimate a set of m=d1m=d-1 directions in parameter space by retraining. Specifically, for each $i=1,\dots,m$, we continue training from $\theta_{w_{k}}$ under a nearby preference $w_{k}^{(i)}$ for $T_{\mathrm{dir}}$ steps to obtain a directional policy $\theta_{w_{k}^{(i)}}$, where $\theta_{w_{k}}$ and $\theta_{w_{k}^{(i)}}$ are all non-dominated. We record $\Delta\theta_{k}^{(i)}=\theta_{w_{k}^{(i)}}-\theta_{w_{k}}$, and $\Delta w_{k}^{(i)}=w_{k}^{(i)}-w_{k}$ as the parameter update vector and the corresponding preference shift. (3) Locally Linear Extension: For each base policy $\theta_{w_{k}}$, we generate candidate policies by extrapolating along the $m$ local directions using $\{\Delta\theta_{k}^{(i)}\}_{i=1}^{m}$. Each candidate is constructed as a linear combination of these directions: $\theta_{k,\mathrm{cand}}=\theta_{w_{k}}+\sum_{i=1}^{m}\alpha_{i}\Delta\theta_{k}^{(i)}$, where step-scale factors $\alpha$ control the step sizes along each direction. We choose each $\alpha_{i}$ from a discrete grid defined by $\alpha_{\mathrm{start}}$, $\alpha_{\mathrm{end}}$, and step size $\Delta\alpha$, resulting in $M$ possible values per direction. Iterating over all combinations across the $m$ directions produces $M^{m}$ candidate policies for each base solution. In parallel, we assign each candidate a matched preference weight $w_{k,\mathrm{cand}}=w_{k}+\sum_{i=1}^{m}\alpha_{i}\,\Delta w_{k}^{(i)}$. (4) Candidate Selection: All candidates $\theta_{k,\mathrm{cand}}$ generated in the locally linear extension stage are evaluated to obtain their respective performance vectors. We then select the subset of non-dominated candidates and forward them to the fine-tuning stage. (5) Preference-Aligned Fine-Tuning: For each selected candidate $\theta$ and its matched weight $w$, we perform a short PPO fine-tuning phase of $T_{\mathrm{ref}}$ steps under $w$ to push the policy closer to the true Pareto front.
Together, these stages explore the local parameter-space structure, enabling efficient tracing of the local Pareto front manifold. This construction yields an interpretable policy representation: local directions in parameter space correspond to meaningful objective trade-offs, allowing parameter changes to be directly associated with movements along the Pareto front.
3.5 Theoretical Analysis
In this section, we present a theoretical analysis for the effectiveness of LLE-MORL, formally modelling the local manifold structure of Pareto-optimal policies and establishing guarantees for the locally linear extension mechanism used to trace the Pareto front. For clarity, we only provide the theory framework and its implications here, while consistent deductions and proofs are located in Appendix B.
First, the continuity of the reconstructed Pareto front by LLE-MORL is guaranteed by the Lipschitz continuity of PPR. Following Theorem 3.2 and Corollary 3.3 provide the theoretical foundation of these properties.
Theorem 3.2.
(Continuity of PPR function $h$.) Suppose that the expected discounted return $V$ is defined by $V:U\rightarrow\mathbb{R}^{d}$, where $U\subseteq\Theta\subseteq\mathbb{R}^{n}$ is an open set. The sufficient condition for Lipschitz continuity of $h(\theta,\Delta\theta)$ with respect to the second variable $\Delta\theta$ is $V\in C^{1}(U)$ and $\forall a\in\mathcal{A},s\in\mathcal{S}$
$\sup_{\theta\in U}\bigl\|\nabla_{\theta}\log\pi_{\theta}(a|s)\bigr\|\leq\mathcal{G}<\infty,$ (2)
The inequality in Theorem 2 above states that the score function $\nabla_{\theta}\log\pi_{\theta}(a|s)$ (Baxter and Bartlett, 2001) is bounded by constant $\mathcal{G}$. This is a common assumption in related research (Aydin and Eksin, 2023; Shi et al., 2024).
Corollary 3.3.
(Surface continuity guaranteed by Lipschitz PPR) Algorithm LLE can construct a continuous surface in the performance space, as an approximation to Pareto front, given appropriate regularity conditions and range of $\alpha$.
Further, we aim to evaluate the error between the reconstructed result by LLE-MORL and the Pareto front $\mathcal{P}$. Assume $\mathcal{P}$ is a differentiable manifold of dimension $m$ embedded in the Euclidean space of dimension $d$, i.e. $\mathcal{P}\in\mathcal{M}^{m}(\mathbb{R}^{d})$, where $\mathrm{dim}\mathcal{P}$ is m=d1m=d-1. Suppose the initialisation returns $\theta^{(0)}$, assuming it corresponds to a non-dominated performance, where $v:=V(\theta^{(0)})\in\mathcal{P}$. Let $P_{v}\subset\mathcal{P}$ be an open neighbourhood of $v$ on the manifold $\mathcal{P}$, where $P_{v}\in\mathcal{M}^{m}(\mathbb{R}^{d})$. Define the neighbour of $v$ to be $P_{v}$, assume there exists a parameter set $\Theta_{v}\subset\Theta$ bijectively mapping to $P_{v}$, denoted by $V|_{\Theta_{v}}:\Theta_{v}\rightarrow P_{v}$.
Theorem 3.4.
(Local lifting manifold) If both $V|_{\Theta_{v}}$ and ${V|_{\Theta_{v}}}^{\!\!-1}$ are $C^{k}$, i.e. $V|_{\Theta_{v}}$ is a diffeomorphism, then $\Theta_{v}$ is also a differentiable manifold embedded in $\mathbb{R}^{n}$ of the same dimension as $P_{v}$, i.e. $\Theta_{v}\in\mathcal{M}^{m}(\mathbb{R}^{n})$.
Therefore, the relation between Pareto front and corresponding parameters can be understand as the diffeomorphism between two differentiable manifold of the same dimension $m$. Next, we define the expression of linear extension.
Condition 3.5.
(Local extension directions) Suppose the directional retrain in LLE-MORL returns $\theta^{(i)}\in\Theta_{v}$, $i=1,\dots,m$. Define
$\mathcal{D}:=\bigl[\theta^{(1)}-\theta^{(0)},\ \dots,\ \theta^{(m)}-\theta^{(0)}\bigr]\in\mathbb{R}^{n\times m}$ (3)
which has full column rank, i.e. $\mathrm{rank}(\mathcal{D})=m$. Further, $\mathrm{span}(\mathcal{D})$ provides an approximation to the tangent space of $\Theta_{v}$ at $\theta^{(0)}$, i.e. $T_{\theta^{(0)}}\Theta_{v}$.
Theorem 3.6.
(Local Pareto error of LLE) Denote the local linear extrapolation by $\tilde{\theta}(\alpha):=\theta^{(0)}+\mathcal{D}\alpha$. Then, for a suitable range of $\alpha\in\mathbb{R}^{m}$, the error of extrapolation to the Pareto front is bounded, with first-order convergence as $\|\alpha\|\rightarrow 0$ and second-order increase as $\|\alpha\|$ grows.
The explicit expression of error bound can be found in Equation 33 in Theorem B.13 in Appendix B. Theorem 3.6 support the effectiveness of LLE-MORL in constructing the Pareto front. And the error is controlled by an appropriate range of $\alpha$, corresponding to the distance of linear extension. The empirical results and discussion is stated in Appendix D.2 for the feasibility of the choice of $\alpha$.
@@ 4 Experiments
(a) MO-Swimmer
(sample-efficient)
(b) MO-Hopper-2d
(sample-efficient)
(c) MO-Ant-2d
(sample-efficient)
(d) MO-Swimmer
(standard-training)
(e) MO-Hopper-2d
(standard-training)
(f) MO-Ant-2d
(standard-training)
Figure 4: Pareto fronts from 2 different settings, comparing our LLE-MORL method with baselines on three 2d continuous-control benchmarks. LLE-MORL consistently achieves wider coverage and closer proximity to the true Pareto front.
4.1 Experiment Setup
In this section, we evaluate the LLE-MORL algorithm using popular continuous MORL benchmark problems from the MO-Gymnasium-v5 (Felten et al., 2023). Our benchmark problems include three two-objective continuous environments: MO-Swimmer, MO-Hopper-2d, MO-Ant-2d, and two three-objective continuous environments: MO-Hopper-3d, MO-Ant-3d. We evaluate the quality of the approximate Pareto front using three standard metrics: Hypervolume (HV), Expected Utility (EU), Sparsity (SP), following the formalism in (Zitzler and Thiele, 2002; Zintgraf et al., 2015; Hayes et al., 2022). Higher HV and EU values indicate better overall front quality, while SP measures the distribution of solutions along the front. Since SP is scale-dependent and less directly related to decision quality, we report it mainly as a complementary metric to assess coverage diversity. More details can be found in Appendix C.
We compare our LLE-MORL against the following state-of-the-art MORL algorithms: (i) GPI-LS (Alegre et al., 2023) applies Generalised Policy Improvement over a discretised set of preference weights and uses linear scalarization to construct a diverse Pareto set. (ii) Concave-augmented Pareto Q-learning (CAPQL) (Lu et al., 2023): learns an ensemble of Q-functions under different preferences and selects actions via conservative aggregation to improve front coverage. (iii) Q-Pensieve (Hung et al., 2023) boosts the sample efficiency of MORL by storing past Q-function snapshots in a replay buffer, enabling explicit policy-level knowledge sharing across training iterations. (iv) MORL/D (Felten et al., 2024) is a deep-RL analogue of decomposition-based multi-objective optimisation that trains subpolicies under scalarised objectives and recombines them via weight decompositions to approximate the Pareto front.
4.2 Results and Analysis on 2D Benchmarks
Table 1: Evaluation of the quality of the Pareto front on 2D-objective benchmarks under different settings.
Setting Environment Metric Method
GPI-LS CAPQL Q-Pensieve MORL/D LLE-MORL-0 LLE-MORL
@@ 4.2 Results and Analysis on 2D Benchmarks
[…]
@@
To assess the performance of MORL, we now present quantitative results evaluating the quality of the approximated Pareto fronts on 2D-objective benchmarks. We conduct experiments under two distinct settings to provide a comprehensive understanding of these algorithms’ capabilities: (1) Sample-Efficient Setting: All methods, including our LLE-MORL approach, were trained under an interaction budget for $1.5\times 10^{5}$ timesteps (including all algorithmic training phases). Given the complexity of continuous control benchmarks, this relatively limited interaction budget serves as a critical testbed for evaluating how rapidly different MORL strategies can discover effective Pareto front approximations. (2) Standard-Training Setting: To assess performance under more common training conditions, most methods, including our LLE-MORL approach, under an interaction budget for $1\times 10^{6}$ timesteps. This setting aligns with common practices for benchmarking in continuous control and allows us to assess the final quality of the Pareto fronts achieved by each algorithm after a more thorough learning process. In all experiments, we fix hyperparameters according to a common setting shared across methods, without assuming any task-specific parameter tuning. Full implementation details are provided in Appendix C.
@@
In the sample-efficient setting, shown in Figure 4 left column and Table 1 top part, LLE-MORL achieves the highest HV and EU in all benchmarks, demonstrating strong capabilities in rapidly achieving high-quality Pareto fronts. Regarding SP, while LLE-MORL does not consistently achieve the leading scores on this metric, its performance generally reflects a good and effective distribution of solutions along the high-quality Pareto fronts it identifies. It should be noted that SP results can be confounded by a fragmentary recovery of the Pareto front. For instance, if Q-Pensieve discovers only two close points of the Pareto front for the MO-Ant problem, the sparsity rating is nearly perfect. Meanwhile, a low SP might also arise from solutions being overly clustered in a small region, as potentially seen with GPI-LS and Q-Pensieve in MO-Swimmer shown in Figure 4(a).
@@
Transitioning to the standard-training setting, shown in Figure 4 right column and Table 1 bottom part, across all benchmarks, LLE-MORL typically achieves the highest HV and highly competitive EU. This superior performance indicates that LLE-MORL finds a more extensive and higher-quality set of solutions, which strongly suggests a better approximation of the Pareto front compared to the baselines. The evidence from Pareto front visualisation further corroborates LLE-MORL’s advantages. In the MO-Swimmer environment, shown in Figure 4(d), LLE-MORL more comprehensively explores the objective space, successfully identifying Pareto optimal solutions in the lower-right region consistently missed by baselines such as GPI-LS, Q-Pensieve and CAPQL. Notably, when comparing the GPI-LS, CAPQL and Q-Pensieve performance to those in the sample-efficient setting for the MO-Swimmer environment, these particular baselines appear to remain constrained by suboptimal solutions in this challenging region, indicating that simply extending training duration did not resolve their exploration deficiencies here. While LLE-MORL’s thorough exploration to achieve this broader coverage means its SP may not be the numerically lowest, the result could be well-justified by the extensive nature of the front.
In summary, LLE-MORL consistently achieves high-quality Pareto fronts across different settings. Its innovative training-free extension step enables efficient generation of diverse solutions, yielding strong performance under limited budgets and consistently improved results with extended training. Runtime comparisons are reported in Appendix C.
4.3 Performance and Scalability in Higher Dimensional Objective Spaces
To evaluate the robustness and discuss the scalability of LLE-MORL beyond standard 2D-objective trade-offs, we extended our experimental evaluation to higher-dimensional scenarios using the MO-Hopper-3d and MO-Ant-3d benchmarks. In these environments, the Pareto front is no longer a one-dimensional curve but a two-dimensional manifold embedded in a three-dimensional performance space. All methods, including our LLE-MORL approach, were trained under an interaction budget for $2\times 10^{6}$ timesteps (including all algorithmic training phases).
Table 2: Evaluation of the quality of the Pareto front on 3D-objective benchmarks.
@@ 4.3 Performance and Scalability in Higher Dimensional Objective Spaces
GPI-LS CAPQL Q-Pensieve MORL/D LLE-MORL-0 LLE-MORL
[…]
The experimental results, reported in Table 2, are highly consistent with the observations from our 2D benchmarks, showing that LLE-MORL maintains a significant advantage in HV and EU over baselines. The robust performance in higher dimensions is a direct consequence of the algorithm’s design, which generalises to $d$-objective tasks without requiring additional algorithmic modifications. As established in our theoretical analysis (Section 3.5), the d1d-1 directional vectors obtained during the retraining phase are sufficient to explore and span the local Pareto manifold. This allows the locally linear extension mechanism to fill in contiguous patches of the Pareto surface effectively.